# Impact of Dataset Properties on Membership Inference Vulnerability of Deep Transfer Learning

### v3

@@ C.4 Proof of Theorem 4
To apply Lemma 2, we will calculate $\mathbb{E}_{z}[q]$ and $\mathbb{E}_{z}[A]$. Let $s_{y_{x}}^{(x)}(x)$ (resp. $s_{y_{x}}^{(z)}(x)$) denote the score of the target $x$ for class $y_{x}$ when the dataset contains $(x,y_{x})$ but not $(z,y_{z})$ (resp. when the dataset contains $(z,y_{z})$ but not $(x,y_{x})$). Let $s_{y_{z}}^{(x)}(z)$ and $z_{y_{z}}^{(z)}(z)$ be corresponding scores of example $z$. Then we have
$s_{y_{x}}^{(x)}(x)$ $=\frac{1}{S}\langle x,\sum_{i=1}^{S}x_{i}+x-x_{S}\rangle$ (A86)
$s_{y_{x}}^{(z)}(x)$ =$\begin{cases}\frac{1}{S}\langle x,\sum_{i=1}^{S}x_{i}+z-x_{S}\rangle\quad&\text{ if }y_{x}=y_{z}\\ \frac{1}{S}\langle x,\sum_{i=1}^{S}x_{i}\rangle\quad&\text{ if }y_{x}\not=y_{z}\end{cases}$ (A87)
$s_{y_{z}}^{(x)}(z)$ =$\begin{cases}\frac{1}{S}\langle z,\sum_{i=1}^{S}z_{i}+x-z_{S}\rangle\quad&\text{ if }y_{x}=y_{z}\\ \frac{1}{S}\langle z,\sum_{i=1}^{S}z_{i}\rangle\quad&\text{ if }y_{x}\not=y_{z}\end{cases}$ (A88)
$s_{y_{z}}^{(z)}(z)$ $=\frac{1}{S}\langle z,\sum_{i=1}^{S}z_{i}+z-x_{S}\rangle$ (A89)
where $x_{i}$ are samples with label $y_{x}$ and $z_{i}$ are samples with label $y_{z}$ when $y_{x}\not=y_{z}$. Thus we obtain
$\mu_{x,x}$ $=\langle x,m_{x}\rangle+\frac{1}{S}\langle x,x-m_{x}\rangle$ (A90)
$\mu_{z,x}$ =$\begin{cases}\langle x,m_{x}\rangle+\frac{1}{S}\langle x,z-m_{x}\rangle\quad&\text{ if }y_{x}=y_{z}\\ \langle x,m_{x}\rangle\quad&\text{ if }y_{x}\not=y_{z}\end{cases}$ (A91)
$\mu_{x,z}$ =$\begin{cases}\langle z,m_{x}\rangle+\frac{1}{S}\langle z,x-m_{x}\rangle\quad&\text{ if }y_{x}=y_{z}\\ \langle z,m_{z}\rangle\quad&\text{ if }y_{x}\not=y_{z}\end{cases}$ (A92)
$\mu_{z,z}$ $=\langle z,m_{x}\rangle+\frac{1}{S}\langle z,z-m_{x}\rangle\quad$ (A93)
$\sigma_{x}$ $=\frac{1}{\sqrt{S}}s||x||$ (A94)
$\sigma_{z}$ $=\frac{1}{\sqrt{S}}s||z||,$ (A95)
where $m_{z}$ is the true class mean of $y_{z}$ (see Appendix B for derivation of $\sigma_{x}$ and $\sigma_{z}$).
@@
$q=$ $\frac{(\mu_{x,x}-\mu_{z,x})(\hat{\mu}_{x,x}-\hat{\mu}_{z,x})}{\hat{\sigma}_{x}^{2}}+\frac{(\mu_{x,z}-\mu_{z,z})(\hat{\mu}_{x,z}-\hat{\mu}_{z,z})}{\hat{\sigma}_{z}^{2}}$ (A96)
$A=$ $\frac{\sigma_{x}}{\hat{\sigma}_{x}^{2}}(\hat{\mu}_{x,x}-\hat{\mu}_{z,x})+\frac{\sigma_{z}}{\hat{\sigma}_{z}^{2}}(\hat{\mu}_{x,z}-\hat{\mu}_{z,z})$ (A97)
In the limit of infinitely many shadow models, estimated statisitcs converge to true parameters, thereby yielding
$q=$ $\left(\frac{\mu_{x,x}-\mu_{z,x}}{\sigma_{x}}\right)^{2}+\left(\frac{\mu_{x,z}-\mu_{z,z}}{\sigma_{z}}\right)^{2}$ (A98)
$A=$ $\frac{\mu_{x,x}-\mu_{z,x}}{\sigma_{x}}+\frac{\mu_{x,z}-\mu_{z,z}}{\sigma_{z}}.$ (A99)
@@
$\mathbb{E}_{z}[|q|]=$ $\Pr_{z}(y_{z}=y_{x})\mathbb{E}_{z}[|q|\mid y_{z}=y_{x}]+\sum_{j=1,j\not=y_{x}}^{C}\Pr_{z}(y_{z}=j)\mathbb{E}_{z}[|q|\mid y_{z}=j]$ (A100)
$=$ $\frac{1}{C}\mathbb{E}_{z}\left[\left.\left(\frac{\langle x,x-z\rangle}{\sqrt{S}s||x||}\right)^{2}+\left(\frac{\langle z,x-z\rangle}{\sqrt{S}s||z||}\right)^{2}\ \right|\ y_{z}=y_{x}\right]$ (A101)
$+\frac{C-1}{C}\mathbb{E}_{z}\left[\left.\left(\frac{\langle x,x-m_{x}\rangle}{\sqrt{S}s||x||}\right)^{2}+\left(\frac{\langle z,z-m_{z}\rangle}{\sqrt{S}s||z||}\right)^{2}\ \right|\ y_{z}\not=y_{x}\right]$ (A102)
$=$ $\frac{1}{CSs^{2}}\mathbb{E}_{z}\left[\left.\frac{\langle x,x-z\rangle^{2}}{||x||^{2}}+\frac{\langle z,x-z\rangle^{2}}{||z||^{2}}\ \right|\ y_{z}=y_{x}\right]$ (A103)
$+\frac{C-1}{CSs^{2}}\mathbb{E}_{z}\left[\left.\frac{\langle x,x-m_{x}\rangle^{2}}{||x||^{2}}+\frac{\langle z,z-m_{z}\rangle^{2}}{||z||^{2}}\ \right|\ y_{z}\not=y_{x}\right],$ (A104)
@@
$\mathbb{E}_{z}[|A|]=$ $\Pr_{z}(y_{z}=y_{x})\mathbb{E}_{z}[|A|\mid y_{z}=y_{x}]+\sum_{j=1,j\not=y_{x}}^{C}\Pr_{z}(y_{z}=j)\mathbb{E}_{z}[|A|\mid y_{z}=j]$ (A105)
$=$ $\frac{1}{C}\mathbb{E}_{z}\left[\left.\left|\frac{\langle x,x-z\rangle}{\sqrt{S}s||x||}+\frac{\langle z,x-z\rangle}{\sqrt{S}s||z||}\right|\ \right|\ y_{z}=y_{x}\right]$ (A106)
$+\frac{C-1}{C}\mathbb{E}_{z}\left[\left.\left|\frac{\langle x,x-m_{x}\rangle}{\sqrt{S}s||x||}+\frac{\langle z,z-m_{z}\rangle}{\sqrt{S}s||z||}\right|\ \right|\ y_{z}\not=y_{x}\right]$ (A107)
$=$ $\frac{1}{C\sqrt{S}s}\mathbb{E}_{z}\left[\left.\left|\frac{\langle x,x-z\rangle}{||x||}+\frac{\langle z,x-z\rangle}{||z||}\right|\ \right|\ y_{z}=y_{x}\right]$ (A108)
$+\frac{C-1}{C\sqrt{S}s}\mathbb{E}_{z}\left[\left.\left|\frac{\langle x,x-m_{x}\rangle}{||x||}+\frac{\langle z,z-m_{z}\rangle}{||z||}\right|\ \right|\ y_{z}\not=y_{x}\right].$ (A109)
Hence we obtain by the Cauchy-Schwarz inequality
$\frac{\mathbb{E}_{z}[|q|]}{\mathbb{E}_{z}[|A|]}=$ $\frac{1}{\sqrt{S}s}\cdot\frac{\mathbb{E}_{z}\left[\frac{\langle x,x-z\rangle^{2}}{||x||^{2}}+\frac{\langle z,x-z\rangle^{2}}{||z||^{2}}\mid y_{z}=y_{x}\right]+(C-1)\mathbb{E}_{z}\left[\frac{\langle x,x-m_{x}\rangle^{2}}{||x||^{2}}+\frac{\langle z,z-m_{z}\rangle^{2}}{||z||^{2}}\mid y_{z}\not=y_{x}\right]}{\mathbb{E}_{z}\left[\left|\frac{\langle x,x-z\rangle}{||x||}+\frac{\langle z,x-z\rangle}{||z||}\right|\mid y_{z}=y_{x}\right]+(C-1)\mathbb{E}_{z}\left[\left|\frac{\langle x,x-m_{x}\rangle}{||x||}+\frac{\langle z,z-m_{z}\rangle}{||z||}\right|\mid y_{z}\not=y_{x}\right]}$ (A110)
$\leq$ $\frac{1}{\sqrt{S}s}\cdot\frac{\mathbb{E}_{z}\left[2||x-z||^{2}\mid y_{z}=y_{x}\right]+(C-1)\mathbb{E}_{z}\left[(||x-m_{x}||^{2}+||z-m_{z}||^{2})\mid y_{z}\not=y_{x}\right]}{\mathbb{E}_{z}\left[2||x-z||\mid y_{z}=y_{x}\right]+(C-1)\mathbb{E}_{z}\left[(||x-m_{x}||+||z-m_{z}||)\mid y_{z}\not=y_{x}\right]}.$ (A111)
@@
$\textsc{tpr}_{\mathrm{RMIA}}(x)$ $\leq 1-F_{|t|}\left(F_{|t|}^{-1}(1-\alpha)-\frac{\mathbb{E}_{z}[|q|]}{\mathbb{E}_{z}[|A|]}\right)$ (A112)
$\leq 1-F_{|t|}\left(F_{|t|}^{-1}(1-\alpha)-\frac{\psi(x,C)}{\sqrt{S}s}\right)$ (A113)
where $F_{|t|}$ is the cdf of the folded normal distribution and
$\psi(x,C)=\frac{\mathbb{E}_{z}\left[2||x-z||^{2}\mid y_{z}=y_{x}\right]+(C-1)\mathbb{E}_{z}\left[(||x-m_{x}||^{2}+||z-m_{z}||^{2})\mid y_{z}\not=y_{x}\right]}{\mathbb{E}_{z}\left[2||x-z||\mid y_{z}=y_{x}\right]+(C-1)\mathbb{E}_{z}\left[(||x-m_{x}||+||z-m_{z}||)\mid y_{z}\not=y_{x}\right]}.$ (A114)
@@
$\textsc{tpr}_{\mathrm{RMIA}}(x)$ $=\Pr_{t}\left(\Pr_{z}(\lambda_{z}+q\geq\log\gamma)\geq\beta\right)$ (A115)
$\leq\Pr_{t}\left(\frac{\mathbb{E}_{z}[|\lambda_{z}|]+\mathbb{E}_{z}[|q|]}{\log(\gamma)}\geq\beta\right)$ (A116)
$=\Pr_{t}\left(|t|\geq F_{|t|}^{-1}(1-\alpha)-\frac{\mathbb{E}_{z}[|q|]}{\mathbb{E}_{z}[|A|]}\right)$ (A117)
$\textsc{fpr}_{\mathrm{RMIA}}(x)$ $=\Pr_{t}\left(\Pr_{z}(\lambda_{z}\geq\log\gamma)\geq\beta\right)$ (A118)
$\leq\Pr_{t}\left(\frac{\mathbb{E}_{z}[|\lambda_{z}|]}{\log(\gamma)}\geq\beta\right)$ (A119)
$=\Pr_{t}\left(|t|\geq F_{|t|}^{-1}(1-\alpha)\right).$ (A120)
We claim that the bound for $\textsc{fpr}_{\mathrm{RMIA}}(x)$ (Equation A119) is tighter than that for $\textsc{tpr}_{\mathrm{RMIA}}(x)$ (Equation A116). Since this is not straightforward, we numerically show this in Table A1. Now, for large enough $S$ we obtain
$\textsc{tpr}_{\mathrm{RMIA}}(x)-\textsc{fpr}_{\mathrm{RMIA}}(x)$ $\leq\Pr_{t}\left(|t|\geq F_{|t|}^{-1}(1-\alpha)-\left|\frac{\mathbb{E}_{z}[q]}{\mathbb{E}_{z}[A]}\right|\right)-\Pr_{t}(|t|\geq F_{|t|}^{-1}(1-\alpha))$ (A121)

### v4

@@ B.3 RMIA power-law
To apply Lemma 5, we will calculate $\mathbb{E}_{\bm{z}}[|q|]$ and $\mathbb{E}_{\bm{z}}[|A|]$. Let $\mathcal{D}_{\mathrm{target}}=\{(\bm{x}_{j,1},j),\dots,(\bm{x}_{j,S},j)\}_{j=1}^{C}$. Let $s_{y_{a}}^{(\bm{b})}(\bm{a})$ denote the score of $(\bm{a},y_{\bm{a}})$ when $\mathcal{D}_{\mathrm{target}}$ contains $(\bm{b},y_{\bm{b}})$ but not the other example. Using similar argument as in the proof of Theorem 2, we have
$s_{y_{\bm{x}}}^{(\bm{x})}(\bm{x})$ $=\frac{1}{S}\langle\bm{x},\sum_{i=1}^{S}\bm{x}_{y_{\bm{x}},i}+\bm{x}-\bm{x}_{y_{\bm{x}},S}\rangle$ (A110)
$s_{y_{\bm{x}}}^{(\bm{z})}(\bm{x})$ =$\begin{cases}\frac{1}{S}\langle\bm{x},\sum_{i=1}^{S}\bm{x}_{y_{\bm{x}},i}+\bm{z}-\bm{x}_{y_{\bm{x}},S}\rangle\quad&\text{ if }y_{\bm{x}}=y_{\bm{z}}\\ \frac{1}{S}\langle\bm{x},\sum_{i=1}^{S}\bm{x}_{y_{\bm{x}},i}\rangle\quad&\text{ if }y_{\bm{x}}\not=y_{\bm{z}}\end{cases}$ (A111)
$s_{y_{\bm{z}}}^{(\bm{x})}(\bm{z})$ =$\begin{cases}\frac{1}{S}\langle\bm{z},\sum_{i=1}^{S}\bm{x}_{y_{\bm{x}},i}+\bm{x}-\bm{x}_{y_{\bm{x}},S}\rangle\quad&\text{ if }y_{\bm{x}}=y_{\bm{z}}\\ \frac{1}{S}\langle\bm{z},\sum_{i=1}^{S}\bm{x}_{y_{\bm{z}},i}\rangle\quad&\text{ if }y_{\bm{x}}\not=y_{\bm{z}}\end{cases}$ (A112)
$s_{y_{\bm{z}}}^{(\bm{z})}(\bm{z})$ $=\frac{1}{S}\langle\bm{z},\sum_{i=1}^{S}\bm{x}_{y_{\bm{z}},i}+\bm{z}-\bm{x}_{y_{\bm{z}},S}\rangle.$ (A113)
Thus we obtain
$\mu_{\bm{x},\bm{x}}$ $=\langle\bm{x},m_{y_{\bm{x}}}\rangle+\frac{1}{S}\langle\bm{x},\bm{x}-m_{y_{\bm{x}}}\rangle$ (A114)
$\mu_{\bm{z},\bm{x}}$ =$\begin{cases}\langle\bm{x},m_{y_{\bm{x}}}\rangle+\frac{1}{S}\langle\bm{x},\bm{z}-m_{y_{\bm{x}}}\rangle\quad&\text{ if }y_{\bm{x}}=y_{\bm{z}}\\ \langle\bm{x},m_{y_{\bm{x}}}\rangle\quad&\text{ if }y_{\bm{x}}\not=y_{\bm{z}}\end{cases}$ (A115)
$\mu_{\bm{x},\bm{z}}$ =$\begin{cases}\langle\bm{z},m_{y_{\bm{x}}}\rangle+\frac{1}{S}\langle\bm{z},\bm{x}-m_{y_{\bm{x}}}\rangle\quad&\text{ if }y_{\bm{x}}=y_{\bm{z}}\\ \langle\bm{z},m_{y_{\bm{z}}}\rangle\quad&\text{ if }y_{\bm{x}}\not=y_{\bm{z}}\end{cases}$ (A116)
$\mu_{\bm{z},\bm{z}}$ $=\langle\bm{z},m_{y_{\bm{x}}}\rangle+\frac{1}{S}\langle\bm{z},\bm{z}-m_{y_{\bm{x}}}\rangle\quad$ (A117)
$\sigma_{\bm{x}}$ $=\frac{1}{\sqrt{S}}\sqrt{\bm{x}^{T}\Sigma\bm{x}}$ (A118)
$\sigma_{\bm{z}}$ $=\frac{1}{\sqrt{S}}\sqrt{\bm{z}^{T}\Sigma\bm{z}},$ (A119)
where $m_{y_{\bm{z}}}$ is the true class mean of class $y_{\bm{z}}$.
@@
$q=$ $\frac{(\mu_{\bm{x},\bm{x}}-\mu_{\bm{z},\bm{x}})(\hat{\mu}_{\bm{x},\bm{x}}-\hat{\mu}_{\bm{z},\bm{x}})}{\hat{\sigma}_{\bm{x}}^{2}}+\frac{(\mu_{\bm{x},\bm{z}}-\mu_{\bm{z},\bm{z}})(\hat{\mu}_{\bm{x},\bm{z}}-\hat{\mu}_{\bm{z},\bm{z}})}{\hat{\sigma}_{\bm{z}}^{2}}$ (A120)
$A=$ $\frac{\sigma_{\bm{x}}}{\hat{\sigma}_{\bm{x}}^{2}}(\hat{\mu}_{\bm{x},\bm{x}}-\hat{\mu}_{\bm{z},\bm{x}})+\frac{\sigma_{\bm{z}}}{\hat{\sigma}_{\bm{z}}^{2}}(\hat{\mu}_{\bm{x},\bm{z}}-\hat{\mu}_{\bm{z},\bm{z}}).$ (A121)
Since $t_{\bm{x}}$ and $t_{\bm{z}}$ follow normal distributions, the location and scale parameters of the true distributions correspond to the mean and standard deviations estimated from infinitely many shadow models, respectively. Thus, we have
$q=$ $\left(\frac{\mu_{\bm{x},\bm{x}}-\mu_{\bm{z},\bm{x}}}{\sigma_{\bm{x}}}\right)^{2}+\left(\frac{\mu_{\bm{x},\bm{z}}-\mu_{\bm{z},\bm{z}}}{\sigma_{\bm{z}}}\right)^{2}$ (A122)
$A=$ $\frac{\mu_{\bm{x},\bm{x}}-\mu_{\bm{z},\bm{x}}}{\sigma_{\bm{x}}}+\frac{\mu_{\bm{x},\bm{z}}-\mu_{\bm{z},\bm{z}}}{\sigma_{\bm{z}}}.$ (A123)
@@
$\mathbb{E}_{\bm{z}}[|q|]=$ $\Pr_{\bm{z}}(y_{\bm{z}}=y_{\bm{x}})\mathbb{E}_{\bm{z}}[|q|\mid y_{\bm{z}}=y_{\bm{x}}]+\sum_{j=1,j\not=y_{\bm{x}}}^{C}\Pr_{\bm{z}}(y_{\bm{z}}=j)\mathbb{E}_{\bm{z}}[|q|\mid y_{\bm{z}}=j]$ (A124)
$=$ $\frac{1}{C}\mathbb{E}_{\bm{z}}\left[\left.\left(\frac{\langle\bm{x},\bm{x}-z\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right)^{2}+\left(\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{\sqrt{S}\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right)^{2}\ \right|\ y_{\bm{z}}=y_{\bm{x}}\right]$ (A125)
$+\frac{C-1}{C}\mathbb{E}_{\bm{z}}\left[\left.\left(\frac{\langle\bm{x},\bm{x}-m_{y_{\bm{x}}}\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right)^{2}+\left(\frac{\langle\bm{z},\bm{z}-m_{y_{\bm{z}}}\rangle}{\sqrt{S}\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right)^{2}\ \right|\ y_{\bm{z}}\not=y_{\bm{x}}\right]$ (A126)
$=$ $\frac{1}{CS}\mathbb{E}_{\bm{z}}\left[\left.\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\ \right|\ y_{\bm{z}}=y_{\bm{x}}\right]$ (A127)
$+\frac{C-1}{CS}\mathbb{E}_{\bm{z}}\left[\left.\frac{\langle\bm{x},\bm{x}-m_{y_{\bm{x}}}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{z}-m_{y_{\bm{z}}}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\ \right|\ y_{\bm{z}}\not=y_{\bm{x}}\right],$ (A128)
@@
$\mathbb{E}_{\bm{z}}[|A|]=$ $\Pr_{\bm{z}}(y_{\bm{z}}=y_{\bm{x}})\mathbb{E}_{\bm{z}}[|A|\mid y_{\bm{z}}=y_{\bm{x}}]+\sum_{j=1,j\not=y_{\bm{x}}}^{C}\Pr_{\bm{z}}(y_{\bm{z}}=j)\mathbb{E}_{\bm{z}}[|A|\mid y_{\bm{z}}=j]$ (A129)
$=$ $\frac{1}{C}\mathbb{E}_{\bm{z}}\left[\left.\left|\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{\sqrt{S}\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\ \right|\ y_{\bm{z}}=y_{\bm{x}}\right]$ (A130)
$+\frac{C-1}{C}\mathbb{E}_{\bm{z}}\left[\left.\left|\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle}{\sqrt{S}\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\ \right|\ y_{\bm{z}}\not=y_{\bm{x}}\right]$ (A131)
$=$ $\frac{1}{C\sqrt{S}}\mathbb{E}_{\bm{z}}\left[\left.\left|\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\ \right|\ y_{\bm{z}}=y_{\bm{x}}\right]$ (A132)
$+\frac{C-1}{C\sqrt{S}}\mathbb{E}_{\bm{z}}\left[\left.\left|\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\ \right|\ y_{\bm{z}}\not=y_{\bm{x}}\right].$ (A133)
Hence we obtain
$\frac{\mathbb{E}_{\bm{z}}[|q|]}{\mathbb{E}_{\bm{z}}[|A|]}\\ =\frac{1}{\sqrt{S}}\cdot\frac{\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}{\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}.$ (A134)
@@
$\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})$ $\leq 1-F_{|Z|}\left(F_{|Z|}^{-1}(1-\alpha)-\frac{\mathbb{E}_{\bm{z}}[|q|]}{\mathbb{E}_{\bm{z}}[|A|]}\right)$ (A135)
$=1-F_{|Z|}\left(F_{|Z|}^{-1}(1-\alpha)-\frac{\Psi(\bm{x},C)}{\sqrt{S}}\right),$ (A136)
where $F_{|Z|}$ is the cdf of the folded normal distribution and
$\Psi(\bm{x},C)=\frac{\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}{\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}.$ (A137)
@@
$\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})$ $=\Pr_{Z}\left(\Pr_{\bm{z}}(\lambda_{\bm{z}}+q\geq\log\gamma)\geq\tau\right)$ (A138)
$\leq\Pr_{Z}\left(\frac{\mathbb{E}_{\bm{z}}[|\lambda_{\bm{z}}|]+\mathbb{E}_{\bm{z}}[|q|]}{\log\gamma}\geq\tau\right)$ (A139)
$=\Pr_{Z}\left(|Z|\geq F_{|Z|}^{-1}(1-\alpha)-\frac{\mathbb{E}_{\bm{z}}[|q|]}{\mathbb{E}_{\bm{z}}[|A|]}\right)$ (A140)
$\textsc{fpr}_{\mathrm{RMIA}}(\bm{x})$ $=\Pr_{Z}\left(\Pr_{\bm{z}}(\lambda_{\bm{z}}\geq\log\gamma)\geq\tau\right)$ (A141)
$\leq\Pr_{Z}\left(\frac{\mathbb{E}_{\bm{z}}[|\lambda_{\bm{z}}|]}{\log\gamma}\geq\tau\right)$ (A142)
$=\Pr_{Z}\left(|Z|\geq F_{|Z|}^{-1}(1-\alpha)\right).$ (A143)
We claim that the bound for $\textsc{fpr}_{\mathrm{RMIA}}(\bm{x})$ (Equation A142) is as tight as that for $\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})$ (Equation A139) for sufficiently large $S$. Let us denote
$\kappa_{\textsc{tpr}}(\gamma)$ $=\Pr_{\bm{z}}(\lambda_{\bm{z}}+q\geq\log\gamma)$ (A144)
$\kappa_{\textsc{fpr}}(\gamma)$ $=\Pr_{\bm{z}}(\lambda_{\bm{z}}\geq\log\gamma).$ (A145)
Since $q=O(1/S)$, for large $S$ we approximate
$\kappa_{\textsc{tpr}}(\gamma)-\kappa_{\textsc{fpr}}(\gamma)\approx p_{\lambda_{\bm{z}}}(\log\gamma)\frac{c_{0}}{S},$ (A146)
for some constant $c_{0}$. Since $\lambda_{\bm{z}}=O(1/\sqrt{S})$, the scaled random variable $\hat{\lambda}_{\bm{z}}=\sqrt{S}\lambda_{\bm{z}}$ is almost independent of $S$. Then by the change of variables formula, we have
$p_{\lambda_{\bm{z}}}(\log\gamma)\frac{c_{0}}{S}=p_{\hat{\lambda}_{\bm{z}}}(\sqrt{S}\log\gamma)\frac{c_{0}}{\sqrt{S}}.$ (A147)
Without loss of generality we may set $\log\gamma=1/\sqrt{S}$. Thus, we have